# f(R) gravity with spacetime torsion

### v1

@@ Abstract
The present work shows the correspondence between $f(R)$ gravity and a dual scalar-tensor theory (with an antisymmetric tensor field) when the affine connection is considered to have an antisymmetric part. It turns out that the $f(R)$ action in presence of spacetime torsion can be recast to a $non-minimally$ coupled scalar-tensor theory with a 2-rank massless antisymmetric tensor field in the Einstein frame, where the scalar field gets coupled with the antisymmetric field through derivative coupling(s).
I Introduction
@@
Despite a huge success of Einstein’s GR starting from the observation of Mercury perihelion to black hole observation along with the detection of gravitational waves, it has various limitations and there have been many thoughts regrading the limitations of GR. From theoretical perspective, one of the main limitations of Einstein’s GR is that the connection (associated with the spacetime metric) is $assumed$ to be symmetric in its lower indices, which in turn leads to a torsion free spacetime. On other hand, the Eisntein-Hilbert action contains only the $linear~{}power$ of Ricci scalar, while the diffeomorphism symmetry demands various combinations of spacetime curvature in the gravitational action. Thus one may argue that the Einstein-Hilbert action does not obey the full diffeomorphism symmetry of the underlying theory. As a result, it turns out that the Einstein’s GR face various challenges during different evolutionary phase of the universe, like the singularity problem (also known as the Big-Bang singularity), the issue of dark matter and dark energy etc.
@@
In order to relax the formal assumption of symmetric connection, the first attempt was made by Cartan to include the torsion in spacetime background, which with further additions was named Einstein-Cartan theory Hehl:1976kj ; Shapiro:2001rz . In particular, it was showed that the spin angular momentum of the matter field(s) can act as a source of spacetime torsion, just like the energy is the source of spacetime curvature. In this regard, there had been a lot of interest to investigate various gravitational aspects of massless Kalb-Ramond field which is a two rank antisymmetric tensor field and has spin one duff ; kalb ; ssg_KR1 ; ssg_KR2 ; ssg_KR3 ; biswarup_KR ; sumanta_KR ; Aashish:2019ykb ; Aashish:2019zsy ; Elizalde:2018rmz ; Elizalde:2018now ; Paul:2020duu ; Paul:2022mup . Regarding the other limitation, it is well known that Einstein–Hilbert action can be generalized by adding higher order curvature terms which naturally arise from diffeomorphism property of the action. Such terms also have their origin in String theory due to quantum corrections. f(R) Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et , Gauss-Bonnet (GB) Nojiri:2005jg ; Nojiri:2005am or more generally Lanczos-Lovelock gravity are some of the candidates in higher curvature gravitational theory. In general inclusion of higher curvature terms in the action leads to the appearance of ghost from higher derivative terms resulting into Ostragradsky instability. The Gauss–Bonnet model (a special case of Lanczos–Lovelock model) is however free of this instability due to appropriate choice of various quadratic combinations of Riemann tensor, Ricci tensor and curvature scalar. In contrast to GB model f(R) gravity model however contains higher curvature terms consisting only of the scalar curvature R. Once again just as GB model, certain classes of f(R) gravity models are free from ghost-like instability. In general f(R) model can be mapped into a scalar–tensor theory at the action level by a conformal transformation of the metric Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et . The issue of instability of the original f(R) model is now reflected in the form of the kinetic and potential terms of the scalar field in the dual scalar–tensor model, where the potential will have a stable minimum and a kinetic term with proper signature. The f(R) gravity theory earned the most attention in the arena of higher curvature gravitational theories, as the f(R) theory can naturally unify the early inflationary phase of the universe with the dark energy era Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et ; Artymowski:2014gea ; Nojiri:2003ft ; Odintsov:2019mlf ; Johnson:2019vwi ; Pinto:2018rfg ; Odintsov:2019evb ; Nojiri:2019riz ; Nojiri:2019fft ; Lobo:2008sg ; Gorbunov:2010bn ; Li:2007xn ; Odintsov:2020nwm ; Odintsov:2020iui ; Appleby:2007vb ; Elizalde:2010ts ; Cognola:2007zu . Various cosmological and astrophysical aspects of higher curvature theories are explored in Nojiri:2010wj ; Nojiri:2017ncd ; Capozziello:2011et ; Artymowski:2014gea ; Nojiri:2003ft ; Odintsov:2019mlf ; Johnson:2019vwi ; Pinto:2018rfg ; Odintsov:2019evb ; Nojiri:2019riz ; Nojiri:2019fft ; Lobo:2008sg ; Gorbunov:2010bn ; Li:2007xn ; Odintsov:2020nwm ; Odintsov:2020iui ; Appleby:2007vb ; Elizalde:2010ts ; Cognola:2007zu ; Li:2007jm ; Carter:2005fu ; Nojiri:2019dwl ; Odintsov:2022unp ; Bamba:2020qdj ; Makarenko:2016jsy ; delaCruzDombriz:2011wn ; Chakraborty:2018scm ; Kanti:2015pda ; Kanti:2015dra ; Saridakis:2017rdo ; Cognola:2006eg .
@@
As mentioned above that without spacetime torsion, f(R) gravity action can be mapped to a minimally coupled scalar-tensor theory in the Einstein frame, where the scalar field appears with a potential that depends on the form of f(R) under consideration. However one may expect that the scenario changes in presence of spacetime torsion, i.e., if the affine connection is considered to have an antisymmetric part. This is the motivation of our present work, in particular, we want to address the correspondence between the Jordan and Einstein frame f(R) gravity in the presence of spacetime torsion. Throughout the paper, a quantity with an overbar represent the same w.r.t. the symmetric Christoffel connection, and a quantity with tilde refers to the Einstein frame quantity.
II Ricci Tensor with torsion
@@
As we know the torsion tensor is the antisymmetric part of the Christoffel connections. We know the Riemann curvature tensor formula is given by
@@
is called the contorsion tensor. The indices are raised and lowered by means of the metric. Here it may be mentioned that the contorsion is antisymmetric in the first two indices, i.e. $K_{abc}=-K_{bac}$, while the torsion tensor $T^{a}_{bc}$ itself is antisymmetric in the last two indices. Due to Eq. (3), the Reimann tensor is written as,
@@
As the covariant derivative can be expanded in the terms of the christoffel connections and its derivatives and then the two terms cancel out giving back the expression found earlier for the ricci tensor $R_{ab}$.Hence, equation (2) gives the expression for the Ricci tensor as a function of the symmetric christoffel connections and the contorsion tensor. This turns out to be sum of the Ricci tensor builded from the symmetric part of the christoffel connections and derivative of contorsion and higher order terms in it. We shall use this expression in the next section to modify the f(R) action in terms of torsion and metric coefficients.
III f(R) action with torsion in Einstein frame
@@
$S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}~{}f\left[g^{bd}(\bar{R}_{bd}-\bar{\nabla}_{a}{K^{a}}_{bd}+\bar{\nabla}_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba})\right]~{}~{}.$ (6)
@@
$\begin{split}\bar{R}=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)+\widetilde{g}^{bd}\Omega^{2}[-\partial_{a}{K^{a}}_{bd}-(\widetilde{\Gamma}^{a}_{ea}+\delta_{a}^{e}\partial_{a}w+\delta_{a}^{a}\partial_{e}w-\widetilde{g}_{ea}\partial^{a}w)K^{e}_{bd}\\ &-(\widetilde{\Gamma}^{e}_{bd}+\delta_{b}^{e}\partial_{d}w+\delta_{d}^{e}\partial_{b}w-\widetilde{g}_{bd}\partial^{e}w){K^{a}}_{ea}+\partial_{d}{K^{a}}_{ba}+(\widetilde{\Gamma}^{a}_{ed}+\delta_{e}^{a}\partial_{d}w+\delta_{d}^{a}\partial_{e}w-\widetilde{g}_{ed}\partial^{a}w){K^{e}}_{ba}\\ &+(\widetilde{\Gamma}^{e}_{ba}+\delta_{b}^{e}\partial_{a}w+\delta_{a}^{e}\partial_{b}w-\widetilde{g}_{ba}\partial^{e}w){K^{a}}_{ed}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}]~{}~{}.\end{split}$ (11)
Actually the first term of Eq.(4) transforms as per Eq.(10), and for other terms, we expand the covariant derivatives in terms of the Christoffel connection and then use the expression of Eq.(9). On simplifying Eq.(11), we get :
$\begin{split}\bar{R}=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)+\Omega^{2}\big{[}2\partial^{e}w{{K^{a}}_{ea}}+\widetilde{g}^{bd}(-\partial_{a}{K^{a}}_{bd}+\partial_{d}{K^{a}}_{ba}\\ &+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}-{\widetilde{\Gamma}^{a}}_{ea}{K^{e}}_{bd}-{\widetilde{\Gamma}^{e}}_{bd}{K^{a}}_{ea}+{\widetilde{\Gamma}^{a}}_{ed}{K^{e}}_{ba}+{\widetilde{\Gamma}^{e}}_{ba}{K^{a}}_{ed})\big{]}~{}~{}.\end{split}$ (12)
@@
$\begin{split}\bar{R}=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)\\ &+\Omega^{2}\big{[}2\partial^{e}w{{K^{a}}_{ea}}+\widetilde{g}^{bd}(-\bar{\nabla}_{a}{K^{a}}_{bd}+\bar{\nabla}_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba})\big{]}~{}~{}.\end{split}$ (13)
@@
$\bar{R}=\Omega^{2}\left[\widetilde{\bar{R}}+6\tilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w-2\partial^{e}w{{K^{a}}_{ea}}+2\bar{\nabla}_{d}{K^{ad}}_{a}+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\right]~{}~{}.$ (14)
Now the 5th term in the above is a total divergent term and by using Gauss divergence theorem we can say that it vanishes at the boundary. Hence the final expression for the relation between $\bar{R}$ and $\widetilde{\bar{R}}$ is:
$\bar{R}=\Omega^{2}\left[\widetilde{\bar{R}}+6\tilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w-2\partial^{e}w{{K^{a}}_{ea}}+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\right]~{}~{}.$ (15)
@@
[…]
$+$ $\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{)}-\Omega^{-4}\left(Af^{\prime}(A)-f(A)\right)\bigg{]}~{}~{}.$
@@
[…]
$+$ $\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{)}-\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}\bigg{]}~{}~{},$
where the suffix ’E’ stands for the Einstein frame, and being a surface term, ~$\widetilde{\Box}w$ vanishes in the Einstein frame action. Eq. (24) can be written in terms of all lowered indices as :
[…]
$+$ $\Omega^{-2}K_{cea}K_{bfd}g^{ca}g^{be}g^{fd}-\Omega^{-2}K_{aed}K_{bfc}g^{ca}g^{be}g^{fd})-\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}\bigg{]}~{}~{}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

# 2A &\&& 2B Raja S.C. Mullick Road Kolkata - 700 032, India

@@ Abstract
The duality between a higher curvature $f(R)$ gravity model and a scalar-tensor theory helps to bring out the role of the additional degree of freedom originating from the higher derivative terms in the gravity action. Such a degree of freedom which appears as a scalar field has been shown to have multiple implications in Cosmological/Astrophysical scenario. The present work proposes a novel generalization to this correspondence between $f(R)$ gravity and a dual scalar-tensor theory when the affine connection is considered to have an antisymmetric part. It turns out that the $f(R)$ action in presence of spacetime torsion can be recast to a $non-minimally$ coupled scalar-tensor theory with a 2-rank massless antisymmetric tensor field in the Einstein frame, where the scalar field gets coupled with the antisymmetric field through derivative coupling(s).
1 Introduction
@@
Despite a huge success of Einstein’s GR starting from the observation of Mercury perihelion to black hole observation along with the detection of gravitational waves, it has various limitations and there have been many modifications beyond Einstein’s gravity to address satisfactory resolution to these limitations. From theoretical perspective, Einstein’s GR begins with the assumption that the connection (associated with the spacetime metric) is symmetric in its lower indices, which in turn leads to a torsion free spacetime. Although so far there is no perceptive experimental signature in favour of space-time torsion , however there is no compelling reason either to assume the absence of non-symmetric part in the affine connection. Moreover the Einstein-Hilbert action contains only the $linear~{}power$ of Ricci scalar, while the diffeomorphism symmetry admits various combinations of spacetime curvature in the gravitational action. Thus one may argue that the Einstein-Hilbert action does not include all allowed terms obeying the full diffeomorphism symmetry of the underlying theory. It is well known that Einstein’s GR face various challenges in explaining various evolutionary phase of the universe, encounters the singularity problem (also known as the Big-Bang singularity) and faces unavoidable divergences in the ultra-violet regime leading to the so called quantum gravity problem.
@@
In order to relax the formal assumption of symmetric connection, the first attempt was made by Cartan to include the torsion in spacetime background, which with further additions was named Einstein-Cartan theory [1, 2]. In particular, it was showed that the spin angular momentum of the matter field(s) can act as a source of spacetime torsion, just like the energy is the source of spacetime curvature. In this regard, there had been a plethora of work to investigate various gravitational aspects of massless Kalb-Ramond field which is a spin one two rank antisymmetric tensor field [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Regarding the other limitation, it is well known that Einstein–Hilbert action can be generalized by adding higher order curvature terms which naturally arise from diffeomorphism property of the action. Such terms also have their origin in String theory due to quantum corrections. $f(R)$ [16, 17, 18], Gauss-Bonnet (GB) [19, 20] or more generally Lanczos-Lovelock gravity are some of the candidates in higher curvature gravitational theory. In general inclusion of higher curvature terms in the action leads to the appearance of ghost from higher derivative terms resulting into Ostragradsky instability. The Gauss–Bonnet model (a special case of Lanczos–Lovelock model) is however free of this instability due to appropriate choice of various quadratic combinations of Riemann tensor, Ricci tensor and curvature scalar. In contrast to GB model $f(R)$ gravity model however contains higher curvature terms consisting only of the scalar curvature R. Once again just as GB model, certain classes of $f(R)$ gravity models are free from ghost-like instability. In general $f(R)$ model can be mapped into a scalar–tensor theory at the action level by a conformal transformation of the metric [16, 17, 18]. The issue of instability of the original $f(R)$ model is now reflected in the form of the kinetic and potential terms of the scalar field in the dual scalar–tensor model, where the potential will have a stable minimum and a kinetic term with proper signature. The $f(R)$ gravity theory earned the most attention in the arena of higher curvature gravitational theories, as the $f(R)$ theory can naturally unify the early inflationary phase of the universe with the dark energy era [16, 17, 18, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Various cosmological and astrophysical aspects of higher curvature theories are explored in [16, 17, 18, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
@@
As mentioned above that without spacetime torsion, $f(R)$ gravity action can be mapped to a minimally coupled scalar-tensor theory in the Einstein frame, where the scalar field appears with a potential that depends on the form of $f(R)$ under consideration. However one may expect that the scenario changes in presence of spacetime torsion, i.e., if the affine connection is considered to have an antisymmetric part. It is therefore crucial to explore whether the torsion included modified $f(R)$ model exhibits a similar dual structure in the form of scalar-tensor theory and how the space-time torsion couples itself with scalar degrees of freedom which originates from the higher derivative part of the action. This is the motivation of our present work, in particular, we want to address the correspondence between the Jordan and Einstein frame $f(R)$ gravity in the presence of spacetime torsion. We also address the issue of Ostragadsky instability in the generalized scenario. Here we would like to mention that an equivalent scalar-tensor representation of Cartan $f(R)$ theory has been demonstrated in [49] where the authors showed that, similar to normal $f(R)$ gravity, Cartan $f(R)$ theory can also be represented to a minimally coupled scalar-tensor theory (without any antisymmetric tensor field). Actually the formalism of the current work is different than [49], in particular, we transform the “$f(R)$ gravity with torsion” from Jordan to Einstein frame based on a conformal transformation of metric, unlike to [49] where the authors rewrite the Cartan $f(R)$ theory in Jordan frame itself by a different fashion without considering any conformal transformation of the metric. This makes our present scenario essentially different from the earlier ones. Throughout the current paper, a quantity with an overbar represent the same with respect to the symmetric Christoffel connection, and a quantity with tilde refers to the Einstein frame quantity.
2 Ricci Tensor with torsion
@@
As we know the torsion tensor is the antisymmetric part of the affine connection. We know the Riemann curvature tensor formula is given by
@@
is called the contorsion tensor. The indices are raised and lowered by means of the metric. Here it may be mentioned that the contorsion is antisymmetric in the first two indices, i.e. $K_{abc}=-K_{bac}$, while the torsion tensor $T^{a}_{bc}$ itself is antisymmetric in the last two indices. Due to Eq. (3), the Riemann tensor is written as,
@@
Owing to the fact that the covariant derivatives (with overbar) can be expanded in the terms of the symmetric Christoffel connection, the above expression gives the expression for the Ricci tensor as a function of the symmetric Christoffel connections and the contorsion tensor. This turns out to be sum of the Ricci tensor built from the symmetric part of the christoffel connections and derivative of contorsion and higher order terms in it. We shall use this expression in the next section to modify the $f(R)$ action in terms of torsion and metric coefficients.
3 $f(R)$ action with torsion in Einstein frame
@@
$\begin{split}S&=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}~{}f\big{[}g^{bd}(\bar{R}_{bd}-\bar{\nabla}_{a}{K^{a}}_{bd}\\ &+\bar{\nabla}_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba})\big{]}~{}~{}.\end{split}$ (6)
@@
$\begin{split}R=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)+\widetilde{g}^{bd}\Omega^{2}[-\partial_{a}{K^{a}}_{bd}\\ &-(\widetilde{\Gamma}^{a}_{ea}+\delta_{a}^{e}\partial_{a}w+\delta_{a}^{a}\partial_{e}w-\widetilde{g}_{ea}\partial^{a}w)K^{e}_{bd}\\ &-(\widetilde{\Gamma}^{e}_{bd}+\delta_{b}^{e}\partial_{d}w+\delta_{d}^{e}\partial_{b}w-\widetilde{g}_{bd}\partial^{e}w){K^{a}}_{ea}\\ &+\partial_{d}{K^{a}}_{ba}+(\widetilde{\Gamma}^{a}_{ed}+\delta_{e}^{a}\partial_{d}w+\delta_{d}^{a}\partial_{e}w-\widetilde{g}_{ed}\partial^{a}w){K^{e}}_{ba}\\ &+(\widetilde{\Gamma}^{e}_{ba}+\delta_{b}^{e}\partial_{a}w+\delta_{a}^{e}\partial_{b}w-\widetilde{g}_{ba}\partial^{e}w){K^{a}}_{ed}\\ &+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}]~{}~{}.\end{split}$ (11)
Actually the first term of Eq.(4) transforms as per Eq.(10), and for other terms, we expand the covariant derivatives (with overbar) in terms of the symmetric Christoffel connection and then use the expression of Eq.(9). On simplifying Eq.(11), we get :
$\begin{split}R=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)+\Omega^{2}\big{[}2\partial^{e}w{{K^{a}}_{ea}}\\ &+\widetilde{g}^{bd}(-\partial_{a}{K^{a}}_{bd}+\partial_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba}\\ &-{\widetilde{\Gamma}^{a}}_{ea}{K^{e}}_{bd}-{\widetilde{\Gamma}^{e}}_{bd}{K^{a}}_{ea}+{\widetilde{\Gamma}^{a}}_{ed}{K^{e}}_{ba}+{\widetilde{\Gamma}^{e}}_{ba}{K^{a}}_{ed})\big{]}~{}~{}.\end{split}$ (12)
@@
$\begin{split}R=&\Omega^{2}(\widetilde{\bar{R}}+6\widetilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w)\\ &+\Omega^{2}\big{[}2\partial^{e}w{{K^{a}}_{ea}}+\widetilde{g}^{bd}(-\bar{\nabla}_{a}{K^{a}}_{bd}\\ &+\bar{\nabla}_{d}{K^{a}}_{ba}+{K^{a}}_{ea}{K^{e}}_{bd}-{K^{a}}_{ed}{K^{e}}_{ba})\big{]}~{}~{}.\end{split}$ (13)
@@
$\begin{split}R&=\Omega^{2}\big{[}\widetilde{\bar{R}}+6\tilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w-2\partial^{e}w{{K^{a}}_{ea}}\\ &+2\bar{\nabla}_{d}{K^{ad}}_{a}+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{]}~{}~{}.\end{split}$ (14)
Now the 5th term in the above is a total divergent term and by using Gauss divergence theorem we can say that it vanishes at the boundary. Hence the final expression for the relation between $R$ and $\widetilde{\bar{R}}$ is:
$\begin{split}R&=\Omega^{2}\big{[}\widetilde{\bar{R}}+6\tilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w-2\partial^{e}w{{K^{a}}_{ea}}\\ &+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{]}~{}~{}.\end{split}$ (15)
@@
$S$ $=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\widetilde{g}}\bigg{[}f^{\prime}(A)\Omega^{-2}\big{(}\widetilde{\bar{R}}+6\tilde{\Box}w-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w$ (16)
-- $2\partial^{e}w{{K^{a}}_{ea}}+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{)}$
-- $\Omega^{-4}\left(Af^{\prime}(A)-f(A)\right)\bigg{]}~{}~{}.$
@@
$S_{\mathrm{E}}=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\widetilde{g}}\bigg{[}\big{(}\widetilde{\bar{R}}-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w-2\partial^{e}w{{K^{a}}_{ea}}$
$+\Omega^{-2}{K^{a}}_{ea}{K^{ed}}_{d}-\Omega^{-2}{K}_{aed}{K}^{eda}\big{)}-\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}\bigg{]}~{}~{},$ (17)
where the suffix ’E’ stands for the Einstein frame, and being a surface term, ~$\widetilde{\Box}w$ vanishes in the Einstein frame action.
Eq. (24) can be written in terms of all lowered indices as :
$S_{\mathrm{E}}$ $=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\widetilde{g}}\bigg{[}\big{(}\widetilde{\bar{R}}-6\widetilde{g}^{\mu\nu}\partial_{\mu}w\partial_{\nu}w$ (18)
-- $2\partial_{b}wK_{cea}\widetilde{g}^{be}g^{ca}+\Omega^{-2}K_{cea}K_{bfd}g^{ca}g^{be}g^{fd}$
-- $\Omega^{-2}K_{aed}K_{bfc}g^{ca}g^{be}g^{fd})$
-- $\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}\bigg{]}~{}~{}$
@@
$\delta S_{\mathrm{E}}$ $=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\widetilde{g}}\bigg{[}-2\partial_{b}w\delta K_{cea}\widetilde{g}^{be}g^{ca}$ (19)
$+$ $\Omega^{-2}(\delta K_{cea}K_{bfd}+K_{cea}\delta K_{bfd})g^{ca}g^{be}g^{fd}$
@@
$\delta S_{\mathrm{E}}$ $=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\widetilde{g}}~{}~{}\delta K_{jkl}\bigg{[}-2\partial^{e}w{\delta^{k}}_{e}g^{jl}$ (20)
$+$ $\Omega^{-2}K_{bfd}g^{jl}g^{bk}g^{fd}+\Omega^{-2}K_{cea}g^{ca}g^{je}g^{kl}$
@@
$S_{\mathrm{E}}$ $=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\widetilde{g}}\bigg{[}\widetilde{\bar{R}}-2\widetilde{g}^{ef}\partial_{e}w\partial_{f}w$ (24)
-- $\Omega^{4}\widetilde{g}^{ca}\widetilde{g}^{be}\widetilde{g}^{fd}\partial_{[e}B_{ad]}\partial_{[f}B_{bc]}$
-- $\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}\bigg{]}~{}~{},$
@@
$S_{\mathrm{E}}$ $=$ $\int d^{4}x\sqrt{-\widetilde{g}}\bigg{[}\frac{\widetilde{\bar{R}}}{2\kappa^{2}}-\frac{1}{2}\widetilde{g}^{ef}\partial_{e}\phi\partial_{f}\phi$ (27)
-- $\frac{Af^{\prime}(A)-f(A)}{f^{\prime}(A)^{2}}-\frac{1}{2}\partial_{[e}Z_{ad]}\partial^{[e}Z^{ad]}$